OpenStudy (jamanorthi):

what is 9a=6 (a+4)??? help plzzz...!!!!!

OpenStudy (anonymous):

First expand the brackets 9a=6a+24 Then collect like terms 9a-6a=24 then simplify 3a=24 and solve for a a=24/3 a=8

OpenStudy (jamanorthi):

but when u substitute 8 for A its not equal

OpenStudy (jamanorthi):

@Mariam.A

OpenStudy (anonymous):

Yes it is, remember your brackets and it will work out nicely as follows 9(8)=6(8)+24 72=48+24 72=72 thus both side are equal

OpenStudy (jamanorthi):

oh okay thank you so much!